OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> 20.3° 10.2° 0.2° 30.3°

OpenStudy (phi):

use u dot v = |u| * |v| * cos(A) A is the angle you want

OpenStudy (anonymous):

42+4*cos(A) 46*cos(A) what is u dot v?

OpenStudy (phi):

u dot v means multiply corresponding numbers and add (x, y) dot (a,b) is x*a+ y*b

OpenStudy (phi):

| (x,y) |^2 = x^2 + y^2 and | (x,y) | = sqrt(x^2 + y^2)

OpenStudy (phi):

so your first step is find the "length" of each vector then the dot product of the vectors. \[ \cos(A) = \frac{ u \cdot v}{|u|\ |v|} \]

OpenStudy (phi):

what is | u | ? u = <6, -1>

OpenStudy (phi):

notice that the length squared can be written as u dot u

OpenStudy (anonymous):

so cos(a)=46/|46|??

OpenStudy (whpalmer4):

A perhaps more intuitive way to go: Find the angle each vector makes with basic trig properties. For example, the vector z = <4,3> has magnitude \[\sqrt{(4-0)^2+(3-0)^2} = 5\] and its angle will be \[\theta = \arctan(3/4) \approx 36.87^\circ\]The difference of the angles will be the angle between them.

OpenStudy (phi):

I would do it step by step first what is | u | ? u = <6, -1>

OpenStudy (phi):

in other words, what is sqrt(6*6 + -1* -1) ?

OpenStudy (anonymous):

sqrt 37

OpenStudy (phi):

yes, now what is the length of v ? |v | = ?

OpenStudy (anonymous):

sqrt 85?

OpenStudy (phi):

use the "sum of squares" then take the square root v = <7, -4> sqrt(7*7 + -4*-4) = sqrt(49+16)= sqrt(65) not sure how you got 85.

OpenStudy (anonymous):

oops i thought it was <7,-6> gotcha

OpenStudy (phi):

notice that v dot v gives you <7, -4> dot <7 , -4>= 7*7 + -4*-4 that is |v|^2 then take the sq root to get the length

OpenStudy (phi):

next, you need u dot v = ?

OpenStudy (phi):

u dot v means multiply corresponding numbers and add (x, y) dot (a,b) is x*a+ y*b

OpenStudy (anonymous):

oh ok one second sorry!

OpenStudy (anonymous):

46?

OpenStudy (phi):

yes, Now we use the fact that the dot product has 2 definitions. There is the multiply corresponding numbers and add version and this one: \[ u \cdot v = |u| \ |v| \ \cos(A) \]

OpenStudy (phi):

If we solve for cos(A), we get \[ \cos(A) = \frac{ u \cdot v}{|u|\ |v|} \]

OpenStudy (anonymous):

so 46/sqrt65sqrt37?

OpenStudy (phi):

what is left is to replace u dot v in the top with 46 and |u| with sqrt(37) and |v| with sq rt(65)

OpenStudy (phi):

time for a calculator.

OpenStudy (phi):

you should get cos(A) = 0.93799...

OpenStudy (anonymous):

i got .938

OpenStudy (phi):

now take the inverse cosine of both sides (in degree mode) A= acos(0,93799)

OpenStudy (anonymous):

oh yeah i got that

OpenStudy (anonymous):

okay

OpenStudy (phi):

and round the answer to the nearest tenth of a degree.

OpenStudy (anonymous):

20.3!

OpenStudy (phi):

success!

OpenStudy (anonymous):

Thank you for being so patient with me and helping me understand! :)

OpenStudy (phi):

whenever you see questions about the angle between vectors, think "dot product"

OpenStudy (anonymous):

okay! :D

OpenStudy (whpalmer4):

As a check: \[|\arctan(\frac{-1}{6}) - \arctan(\frac{-4}{7})| = |-29.74^\circ-(-9.462^\circ)| \approx 20.3^\circ\]